OpenStudy (anonymous):

PROBABILITY DISTRIBUTION AND EXPECTED VALUE QUESTION An online gambling site offers a first prize of $50,000 and two second prizes of $10,000 each for registered users when they place a bet. A random bet will be selected over a 24-hour period. One million bets are received in the contest. Find the expected winnings if you can place one registered bet of $1 in the given period. @satellite73

OpenStudy (anonymous):

this is easy to do if we think this way buy up all the tickets, you spend \(\$1,000,000\)

OpenStudy (anonymous):

you get back a total of \(\$500,000+2\times \$10,000=\$520,000\)

OpenStudy (anonymous):

it is 50000 not 500000

OpenStudy (anonymous):

oh even worse

OpenStudy (anonymous):

you get a back a total of \(50,000+20,000=70,000\) for a net loss of \[1,000,000-70,000= -930000\]

OpenStudy (anonymous):

averaged over the one million tickets your loss per ticket is \[-.93\]

OpenStudy (anonymous):

ok.sure we not done right.lol

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no we are done

OpenStudy (anonymous):

your expected value is \(-.93\) i.e. you expect to lose 93 cents for every one dollar ticket you buy

OpenStudy (anonymous):

so u divided the -9300000 by 1000000 to get -.93 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well actually it is \(-930,000\) divided by \(1,000,000\)

OpenStudy (tkhunny):

Or, for n = 1,000,000, we have \(\left(\dfrac{1}{n}\cdot 50000\right) + \left(\dfrac{2}{n}\cdot 10000\right) + \left(\dfrac{n-3}{n}\cdot 0\right) - 1 = 0.93\), wonderfully the same correct solution.

OpenStudy (tkhunny):

Did it again! -0.93!

OpenStudy (anonymous):

are ur last figures in parentheses multiplied by 0

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

or even \[49,999\times \frac{1}{1,000,000}+9,999\times \frac{2}{1,000,000}-1\times \frac{999,9997}{1,000,000}\]

OpenStudy (anonymous):

but i always think it is easier to do it the first way less confusing for me

OpenStudy (tkhunny):

Yes, multiplied by zero. This is just the piece that represents winning nothing. You just gave them your dollar! Obviously, this term could have been omitted.

OpenStudy (anonymous):

or "the sum of the amount you win times the probability you win it"

OpenStudy (tkhunny):

Thoroughly agree @satellite73 !! There are several ways to proceed. The fundamental principles are the same. For a learning exercise, I would ALWAYS try more than one way to achieve the same answer.

OpenStudy (anonymous):

so the -1 represents the $1 right ?

OpenStudy (tkhunny):

Right. This is particularly clear in @satellite73's last version. 49999 = 50000 - 1 9999 = 10000 - 1 There's that dollar that you have to pay just to get in the game.

OpenStudy (anonymous):

perfect